# A variational approach to dimension-free self-normalized concentration

### v1

@@ Abstract
We study the self-normalized concentration of vector-valued stochastic processes. We focus on bounds for sub-$\psi$ processes, a tail condition that encompasses a wide variety of well-known distributions (including sub-exponential, sub-Gaussian, sub-gamma, and sub-Poisson distributions). Our results recover and generalize the influential bound of Abbasi-Yadkori et al. [1] and fill a gap in the literature between determinant-based bounds and those based on condition numbers. As applications we prove a Bernstein inequality for random vectors satisfying a moment condition (which is more general than boundedness), and also provide the first dimension-free, self-normalized empirical Bernstein inequality. Our techniques are based on the variational (PAC-Bayes) approach to concentration.
@@ 1 Introduction
What about more general classes of processes? Recently, Whitehouse et al. [60] gave self-normalized concentration bounds for vector-valued sub-$\psi$ processes, a powerful and general abstraction first introduced by Howard et al. [33] which captures many distributions of interest, including sub-Poisson, sub-exponential, sub-gamma, among others. While comprehensive, the bounds of Whitehouse et al. [60] bounds deviate from the aforementioned log-determinant bounds in two ways. First, they are dimension-dependent and second, they scale with the condition number of $V_{t}$ instead of the determinant. Neither dependency is strictly better than the other; both can be tighter in different regimes. However, the discrepancy invites the following question:
@@
Our techniques are based on the variational approach to concentration (sometimes called the PAC-Bayesian approach [10, 11]), which has been successfully leveraged in recent years to study the concentration of random vectors and matrices [63, 49, 48, 65, 14]. As we will explain below, the variational approach is comparable to a data-dependent method of mixtures, in which we choose the mixture distribution to be data-dependent but pay the price of this dependency with an $f$-divergence.
While closing the gap between determinant-based and condition number-based bounds is mathematically interesting, it is also practically relevant. Beyond their application in bandit problems [1, 2, 56, 12, 39, 59], self-normalized bounds are used in many areas, such as system identification [44], control of dynamic systems [38], estimation in autoregressive processes and diffusion-limited aggregation processes [6], sequential inference in time-series [53], and Markov decision-processes [62, 56]. Moreover, in many of these applications, matrices are often ill-conditioned due to numerical instability or anisotropy [31]. Determinant-based bounds which omit the dependence on condition numbers can thus improve results in these areas.
@@ 1.1 Related Work
De la Peña and others then built off of this work to provide the foundations of a general theory of self-normalized concentration, best summarized in the 2008 book of de la Peña, Lai, and Shao [19]. In the scalar setting, subsequent work by Bercu and Touati [5, 6] extended these results to heavy-tailed increments, developing self-normalized bounds when the normalization process $(V_{t})$ is the predictable or total quadratic variation. In 2020, Howard et al. [33] further generalized this framework by the studying time-uniform concentration of sub-$\psi$ processes—processes characterized by specific tail conditions on the cumulant generating functions—thereby providing a general, unifying framework for both self-normalized and non-self-normalized bounds.
Beyond the scalar setting, de la Peña et al. [16] (see also de la Peña et al. [19, Chapter 14]) considered the self-normalized concentration of vector-valued processes, which take the form of bounds on $\|S_{t}\|_{V_{t}^{-1}}$ for a vector-valued process $(S_{t})$ and a matrix-valued process $(V_{t})$. Hotelling’s $T^{2}$-statistic [32]—the generalization of the Student t-statistic to the multivariate setting—is an example of such a process. De la Peña [16] give bounds on the probability that $\|S_{t}\|_{V_{t}^{-1}}$ belongs to a convex set, when $S$ and $V$ obey a “generalized canonical” assumption (see [19, Equation (14.5)]). This assumption is similar to, but weaker than, the sub-$\psi$ assumption of Howard et al. [33] (extended to $\mathbb{R}^{d}$) that we employ (Definition 2.1). Unfortunately, outside of the sub-Gaussian case, explicit expressions for these sets appear unavailable.
In the contextual bandit setting with sub-Gaussian noise, Abbasi-Yadkori et al. [1] apply results of de la Peña et al. [18] and give what is now a famous self-normalized inequality. They show that, if $S_{t}=\sum_{k\leq t}X_{k}\epsilon_{k}$ and $V_{t}=U_{0}+\sum_{k\leq t}X_{k}X_{k}^{\intercal}$ where $\epsilon_{k}$ is scalar sub-Gaussian and $X_{k}\in\mathbb{R}^{d}$ is predictable (i.e., measurable at time k1k-1) then $\|S_{\tau}\|_{V_{\tau}^{-1}}\lesssim\sqrt{\log(\det V_{\tau})}$ at all stopping times $\tau$. This bound was extended to infinite-dimensional spaces by Chowdhury and Gopalan [12].
Recently, Whitehouse et al. [60] extended the sub-$\psi$ condition of Howard et al. [33] to $\mathbb{R}^{d}$ and showed that, if $(S_{t})$ and $(V_{t})$ obey such a condition, then
@@
Neither $\det V_{\tau}$ (the product of all eigenvalues) or $\kappa(V_{\tau})$ dominates the other, making the bound of Whitehouse et al. [60] and Abbasi-Yadkori et al. [1] incomparable in general. If the matrices are well-conditioned then the former may be tighter, depending on the dimension. If there are many small eigenvalues (e.g., perhaps $V_{t}$ is the result of many rank $k$ updates for $k\ll d$) then the determinant-based bound may be tighter. In any case, the gap between condition number-based bounds and determinant-based bounds is what motivates this work.
Very recently, both Akhavan et al. [3] Metelli et al. [46] gave dimension-free, self-normalized Bernstein inequalities, building off the earlier work of Faury et al. [22] and Zhou et al. [64], both of which provide dimension-dependent Bernstein-type bounds. Akhavan et al. [3] study the case in which the normalization process $(V_{t})$ is the quadratic variation of $(S_{t})$ and Metelli et al. [46] study the bandit setting (as do Faury et al. [22] and Zhou et al. [64]). All of these results apply to bounded processes and do not extend to general sub-$\psi$ processes. They are closest to the Bennett and Bernstein results we give in Section 5, though we note that our Bernstein bound applies to random vectors satisfying a more general condition than boundedness. See Table 1 for an overview of known self-normalized concentration results in the multivariate setting.
@@
Chowdhury and Gopalan [12] sub-Gaussian $\det V_{t}$
@@
Table 1: A brief overview of previous vector-valued self-normalized concentration inequalities. “Condition” refers to the distributional assumptions made on the underlying process. “UB on $\|S_{t}\|$” means the bound is a function of the upper bound of the norm of the observations (i.e., it is not adaptive to $V_{t}$). We note that Kirschner et al. [36] provide two bounds: one under a Gaussian assumption and one under a bounded assumption. “Gen. canonical” refers to the generalized canonical assumption of de la Peña et al. [19, Section 14.1.2]. () indicates that the results are not in closed-form. (∗∗) indicates that these works were concurrent to this one.
Finally, let us discuss previous work on the variational approach to concentration, which undergirds our work here. This technique originated in PAC-Bayesian approach to statistical learning theory as a method to bound the risk of randomized predictors. As far as we are aware, the first to employ the variational approach explicitly for the purposes of multivariate concentration was Oliveira [49] in 2016, who was interested in bounding the tails of quadratic forms. Since then, a number of authors of have employed it to various problems.
In 2018, Giulini [27] used the variational approach to estimate the Gram operator in Hilbert spaces. Around the same time, Catoni and Giulini [10, 11] used it to estimate the mean of random vectors and the operator norm of random matrices in finite-dimensional Euclidean spaces. More recently, both Nakakita et al. [48] and Zhivotovskiy [63] return to the question of matrix concentration, each giving dimension-free bounds on sums of random matrices under various conditions. Chugg et al. [14] study the concentration of random vectors using the variational method, giving bounds under a variety of distributional assumptions (including sub-$\psi$ condition which we study here).
All of the above works using the variational approach are in non-self-normalized settings. By contrast, Kirschner et al. [36] and Ziemann [65] recently noted that variational ideas can be used to obtain bounds similar to (or identical to) the ones presented by Abbasi-Yadkori et al. [1]. Ziemann [65] also gives a (dimension-dependent) Bernstein inequality for bounded observations in a contextual bandit setting. Overall, one can view our work here as extending the ideas of Kirschner et al. [36] and Ziemann [65] to more general stochastic processes.
@@ 1.2 Contributions and outline
[…]
Theorems 4.3 and 4.7 employ a method known as stitching [34] which iteratively applies Theorem 4.1 with different parameters in order to give a bound which remains tight at all times. This procedure costs only an iterated logarithm term. Theorem 4.3 applies this technique to sub-gamma processes specfically and obtains precise constants, while Theorem 4.7 applies to more general processes but is asymptotic only.
@@ 2 Preliminaries
Fix a filtered probability space $(\Omega,\mathcal{A},\mathcal{F}\equiv(\mathcal{F}_{t})_{t\geq 0},P)$. We assume that we are working in discrete time, so $\mathcal{F}$ is indexed by $\mathbb{N}\cup\{0\}$. [^1] Though we expect that our techniques are general enough to extend to continuous time as well. We consider two stochastic processes $(S_{t})_{t\geq 1}$ and $(V_{t})_{t\geq 0}$ on $(\Omega,\mathcal{A},\mathcal{F},P)$ which are adapted to $\mathcal{F}$ (i.e., $S_{t}$ and $V_{t}$ are $\mathcal{F}_{t}$-measurable). Throughout this work we assume that $S_{t}\in\mathbb{R}^{d}$ and $V_{t}\in\mathbb{R}^{d\times d}$ for some finite positive integer $d$. We assume that $V_{t}$ is positive-definite for each $t\geq 0$, meaning that $\langle\theta,V_{t}\theta\rangle\geq 0$ for all $\theta\in\mathbb{R}^{d}$.
We are interested obtaining bounds on $\|S_{\tau}\|_{V_{\tau}^{-1}}$ for any $\mathcal{F}$-adapted stopping time $\tau$ under various assumptions on the relationship between $S_{t}$ and $V_{t}$. More specifically, we think of $(V_{t})$ as the accumulated variance process of $(S_{t})$ or, in the words of Blackwell and Freedman [7], a measure of the intrinsic time-scale of $(S_{t})$. This is formalized with the notion of a sub-$\psi$ process, which generalizes many distributions of interest. Let $\mathbb{S}^{d-1}:=\{x\in\mathbb{R}^{d}:\|x\|_{2}=1\}$ denote the unit sphere in $\mathbb{R}^{d}$. Recall that a process $(A_{t})$ is a supermartingale with respect to $\mathcal{F}$ if it is $\mathcal{F}$-adapted and $\mathbb{E}[A_{t}|\mathcal{F}_{t-1}]\leq A_{t-1}$ for all $t\geq 1$. [^2] One would typically specify the distribution $P$ under which $(A_{t})$ is a supermartingale, and refer to $(A_{t})$ as a $P$-supermartingale. However, in our case we consider only one probability measure $P$.
@@
Let $\psi:[0,\lambda_{\max})\to\mathbb{R}_{\geq 0}$. Let $(S_{t})$ be an $\mathbb{R}^{d}$-valued process and $(V_{t})$, $V_{t}\succ 0$ an $\mathbb{R}^{d\times d}$-valued process, both adapted to a filtration $\mathcal{F}$. We call $(S_{t})$ a sub-$\psi$ process with variance proxy $(V_{t})$ if, for all $\theta\in\mathbb{S}^{d-1}$, all $\lambda\in[0,\lambda_{\max})$, and all $t\geq 0$,
@@
where $(L_{t}(\theta))_{t\geq 0}$ is a nonnegative supermartingale adapted to $(\mathcal{F}_{t})$. Equivalently, we say that $(S_{t},V_{t})$ is a sub-$\psi$ process.
For scalar $S_{t}$ and $V_{t}$, sub-$\psi$ processes were originally proposed and studied by Howard et al. [33], building off of earlier work of Freedman [24] and de la Peña et al. [18]. The definition of a sub-$\psi$ process was then generalized to $\mathbb{R}^{d}$ by Whitehouse et al. [60]. Like both Howard et al. [33] and Whitehouse et al. [60], we assume that $\psi$ is CGF-like, meaning that it is strictly convex and twice continuously differentiable, with $\psi(0)=\lim_{\lambda\to 0^{+}}\textnormal{d}\psi(\lambda)/\textnormal{d}\lambda=0$.
While the definition of a sub-$\psi$ process may appear abstract at first glance, let us reassure the reader that the definition captures many familiar examples.
@@
    $\psi_{N}=\psi_{N}(\lambda):=\lambda^{2}/2$ for $\lambda\in[0,\lambda_{\max})$ results in a sub-exponential process distribution (sub-Gaussian if $\lambda_{\max}=\infty)$. If $X_{t}$ is $\Sigma_{t}$-sub-Gaussian [^3] That is, $\mathbb{E}[\exp(\langle\theta,X_{t}\rangle)|\mathcal{F}_{t-1}]\leq\exp(\langle\theta,\Sigma_{t}\theta\rangle]$ for all $\theta\in\mathbb{S}^{d-1}$. then $S_{t}=\sum_{k\leq t}X_{k}$ is a sub-$\psi_{N}$ process with $\lambda_{\max}=\infty$ and $V_{t}=\sum_{k\leq t}\Sigma_{k}$. More exotically, consider any sequence of random vectors $(X_{t})$ that are conditionally symmetric: $X_{t}\sim-X_{t}|\mathcal{F}_{t-1}$ for all $t$. Lemma 3 of de la Peña et al. [15] implies that $S_{t}=\sum_{k\leq t}X_{k}$ and $V_{t}=\sum_{k\leq t}X_{k}X_{k}^{\intercal}$ define a sub-$\psi_{N}$ process with $\lambda_{\max}=\infty$. As in the scalar case, bounded random vectors can also be shown to be sub-$\psi_{N}$.
@@
[…]
@@
[…]
@@
[…]
@@
To prove our results, we instead turn to a related but distinct technique which we call the variational approach to concentration. The variational approach also begins with a mixture but after step (2) applies the Donsker-Varadhan change of measure formula in order to bound $Z_{t}$ in terms of the KL divergence between $\rho$ and some prior distribution. The benefit of this technique is that $\rho$ can be data-dependent. This approach lies at the heart of (and originated in) PAC-Bayesian learning theory [30, 4]. A generic template for such a bound, which applies to general stochastic processes and handles stopping times, was given by Chugg et al. [13, Theorem 4]. See also Flynn et al. [23] for a similar statement.

### v2

@@ Abstract
We study the self-normalized concentration of vector-valued stochastic processes. We focus on bounds for “sub-$\psi$” processes, a well-known and quite general class of processes that encompasses a wide variety of well-known tail conditions (including sub-exponential, sub-Gaussian, sub-gamma, sub-Poisson, and several heavy-tailed settings without a moment generating function such as symmetric or bounded 2nd or 3rd moments). Our results recover and generalize the influential bound of de la Peña et al. [17] (proved again in Abbasi-Yadkori et al. [1]) in the sub-Gaussian case. Further, we fill a gap in the literature between determinant-based bounds and more recent bounds based on condition numbers. As applications we prove a Bernstein inequality for random vectors satisfying a moment condition (a more general condition than boundedness), and also provide the first dimension-free self-normalized empirical Bernstein inequality. Our techniques are based on the variational (PAC-Bayes) approach to concentration.
Contents
@@ 1 Introduction
What about more general classes of processes? Recently, Whitehouse et al. [63] gave self-normalized concentration bounds for vector-valued sub-$\psi$ processes, a powerful and general abstraction first introduced by Howard et al. [34] which captures many distributions of interest, including sub-Poisson, sub-exponential, sub-gamma, in addition to several heavy tailed settings where the moment generating function may not exist (see Section 2). While comprehensive, the bounds of Whitehouse et al. [63] bounds deviate from the aforementioned log-determinant bounds in two ways. First, they are dimension-dependent and second, they scale with the condition number of $V_{t}$ instead of the determinant. Neither dependency is strictly better than the other; both can be tighter in different regimes. However, the discrepancy invites the following question:
@@
Our techniques are based on the variational approach to concentration (sometimes called the PAC-Bayesian approach [11, 12]), which has been successfully leveraged in recent years to study the concentration of random vectors and matrices [65, 51, 50, 67, 16]. The variational approach is comparable to a data-dependent method of mixtures, in which we choose the mixture distribution to be data-dependent but pay the price of this dependency with an $f$-divergence.
While closing the gap between determinant-based and condition number-based bounds is mathematically interesting, it is also practically relevant. Beyond their application in bandit problems [1, 2, 58, 13, 40, 62], self-normalized bounds are used in many areas, such as system identification [46], control of dynamic systems [39], estimation in autoregressive processes and diffusion-limited aggregation processes [7], sequential inference in time-series [55], and Markov decision-processes [64, 58]. Moreover, in many of these applications, matrices are often ill-conditioned due to numerical instability or anisotropy [32]. Determinant-based bounds which omit the dependence on condition numbers can influence results in these areas.
@@ 1.1 Related Work
De la Peña and others then built off of this work to provide the foundations of a general theory of self-normalized concentration, best summarized in the 2008 book of de la Peña, Lai, and Shao [20]; the book is mostly focused on the scalar setting (with some sub-Gaussian vector bounds as we describe in the following paragraph). In the scalar setting, subsequent work by Bercu and Touati [6, 7] extended these results to heavy-tailed increments, developing self-normalized bounds when the normalization process $(V_{t})$ is the predictable or total quadratic variation. In 2020, Howard et al. [34] further generalized this framework by the studying time-uniform concentration of sub-$\psi$ processes—processes characterized by specific tail conditions on the cumulant generating functions—thereby providing a general, unifying framework for both self-normalized and non-self-normalized scalar (and matrix) bounds.
Beyond the scalar setting, de la Peña et al. [18] (see also de la Peña et al. [20, Chapter 14]) considered the self-normalized concentration of vector-valued processes, which take the form of bounds on $\|S_{t}\|_{V_{t}^{-1}}$ for a vector-valued process $(S_{t})$ and a matrix-valued process $(V_{t})$. Hotelling’s $T^{2}$-statistic [33]—the generalization of the Student t-statistic to the multivariate setting—is an example of such a process. De la Peña [18] give bounds on the probability that $\|S_{t}\|_{V_{t}^{-1}}$ belongs to a convex set, when $S$ and $V$ obey a “generalized canonical” assumption (see [20, Equation (14.5)]). This assumption is similar in spirit to the sub-$\psi$ assumption of Howard et al. [34] (extended to $\mathbb{R}^{d}$) that we employ (Definition 2.1). Unfortunately, outside of the sub-Gaussian case, explicit expressions for these sets appear unavailable.
In the contextual bandit setting with sub-Gaussian noise, Abbasi-Yadkori et al. [1] apply results of de la Peña et al. [17] and give what is now a famous self-normalized inequality. They show that, if $S_{t}=\sum_{k\leq t}X_{k}\epsilon_{k}$ and $V_{t}=U_{0}+\sum_{k\leq t}X_{k}X_{k}^{\intercal}$ where $\epsilon_{k}$ is scalar sub-Gaussian and $X_{k}\in\mathbb{R}^{d}$ is predictable (i.e., measurable at time k1k-1) then $\|S_{\tau}\|_{V_{\tau}^{-1}}\lesssim\sqrt{\log(\det V_{\tau})}$ at all stopping times $\tau$. This bound was extended to the infinite-dimensional case in Abbasi-Yadkori [3].
Recently, Whitehouse et al. [63] extended the general sub-$\psi$ condition of Howard et al. [34] to $\mathbb{R}^{d}$ and showed that, if $(S_{t})$ and $(V_{t})$ obey such a condition, then
@@
This qualitatively new bound was also shown to be sharp with a corresponding lower bound. Neither $\det V_{\tau}$ (the product of all eigenvalues) or $\kappa(V_{\tau})$ dominates the other, making the bound of Whitehouse et al. [63] and Abbasi-Yadkori et al. [1] incomparable in general. If the matrices are well-conditioned then the former may be tighter, depending on the dimension. If there are many small eigenvalues (e.g., perhaps $V_{t}$ is the result of many rank $k$ updates for $k\ll d$) then the determinant-based bound may be tighter. In any case, the gap between condition number-based bounds and determinant-based bounds is what motivates this work.
Very recently, both Akhavan et al. [4] Metelli et al. [48] gave dimension-free, self-normalized Bernstein inequalities, building off the earlier work of Faury et al. [24] and Zhou et al. [66], both of which provide dimension-dependent Bernstein-type bounds. Akhavan et al. [4] study the case in which the normalization process $(V_{t})$ is the quadratic variation of $(S_{t})$ and Metelli et al. [48] study the bandit setting (as do Faury et al. [24] and Zhou et al. [66]). All of these results apply to bounded processes and do not extend to general sub-$\psi$ processes. They are closest to the Bennett and Bernstein results we give in Section 5, though we note that our Bernstein bound applies to random vectors satisfying the Bernstein condition—a more general condition than boundedness.
@@
Abbasi-Yadkori [3] sub-Gaussian $\det V_{t}$
@@
Table 1: A brief overview of previous vector-valued self-normalized concentration inequalities. “Condition” refers to the distributional assumptions made on the underlying process. “UB on $\|S_{t}\|$” means the bound is a function of the upper bound of the norm of the observations (i.e., it is not adaptive to $V_{t}$). We note that Kirschner et al. [37] provide two bounds: one under a Gaussian assumption and one under a bounded assumption. “Gen. canonical” refers to the generalized canonical assumption of de la Peña et al. [20, Section 14.1.2]. The difference between Abbasi-Yadkori et al. [1] and Abbasi-Yadkori [3] is that the latter holds in infinite-dimensional spaces. () indicates that the results are not in closed-form. (∗∗) indicates that these works were concurrent to this one.
In the contextual bandit setting, Martinez-Taboada et al. [43] also give a dimension-free determinant-based bound for random vectors satisfying the Bernstein condition. It differs from ours in several ways: It is time-uniform for the finite horizon $1\leq t\leq n$ for some fixed $n$, it has a multiplicative dependence between $\log(1/\delta)$ and $\log\det(V_{t})$, and it requires a deterministic upper bound on $\log\det(V_{t})$ to instantiate. Nevertheless, it has several appealing properties, and will dominate our bound in some regimes. We provide a more involved discussion in Section 4.3. See Table 1 for an overview of known self-normalized concentration results in the multivariate setting.
Finally, let us discuss previous work on the variational approach to concentration, which undergirds our work here. This technique originated in PAC-Bayesian approach to statistical learning theory as a method to bound the risk of randomized predictors. As far as we are aware, the first to employ the variational approach explicitly for the purposes of multivariate concentration was Oliveira [51] in 2016, who was interested in bounding the tails of quadratic forms. Since then, a number of authors of have employed it to solve various problems.
In 2018, Giulini [28] used the variational approach to estimate the Gram operator in Hilbert spaces. Around the same time, Catoni and Giulini [11, 12] used it to estimate the mean of random vectors and the operator norm of random matrices in finite-dimensional Euclidean spaces. More recently, both Nakakita et al. [50] and Zhivotovskiy [65] return to the question of matrix concentration, each giving dimension-free bounds on sums of random matrices under various conditions. Chugg et al. [16] study the concentration of random vectors using the variational method, giving bounds under a variety of distributional assumptions (including sub-$\psi$ condition which we study here). All of the above works using the variational approach are in non-self-normalized settings.
Kirschner et al. [37] and Ziemann [67] recently noted that variational ideas can be used to obtain bounds similar to (or identical to) the ones presented by Abbasi-Yadkori et al. [1]. Ziemann [67] also gives a dimension-dependent Bernstein inequality for bounded observations in a contextual bandit setting. One can view our work here as extending the ideas of Kirschner et al. [37] and Ziemann [67] to much more general stochastic processes.
@@ 1.2 Contributions and outline
Section 4 then provides concentration results for general sub-$\psi$ processes evolving in $\mathbb{R}^{d}$. In particular, Theorem 4.1 provides a dimension-free, self-normalized “line-crossing inequality,” which gives the probability that $\|S_{t}\|_{V_{t}^{-1}}$ ever crosses a threshold which depends on a parameter $\lambda>0$. Different values of $\lambda$ optimize the threshold at different (intrinsic) times.
Theorems 4.4 and 4.9 employ a method known as stitching [35] which iteratively applies Theorem 4.1 with different parameters in different geometrically-spaced epochs of time, in order to give a bound which remains tight at all times. This procedure costs only an iterated logarithm term (this is similar to a doubling trick, chaining, peeling, etc.). Theorem 4.4 applies this technique to sub-gamma processes specifically and obtains precise constants, while Theorem 4.9 applies to more general processes but is asymptotic only.
@@ 2 Preliminaries
Fix a filtered probability space $(\Omega,\mathcal{F}\equiv(\mathcal{F}_{t})_{t\geq 0},P)$. We assume that we are working in discrete time, so $\mathcal{F}$ is indexed by $\mathbb{N}\cup\{0\}$. We consider two stochastic processes $(S_{t})_{t\geq 1}$ and $(V_{t})_{t\geq 0}$ on $(\Omega,\mathcal{F},P)$ which are adapted to $\mathcal{F}$ (i.e., $S_{t}$ and $V_{t}$ are $\mathcal{F}_{t}$-measurable). Throughout this work we assume that $S_{t}\in\mathbb{R}^{d}$ and $V_{t}\in\mathbb{R}^{d\times d}$ for some finite positive integer $d$; $S$ is a mnemonic for sum, and $V$ for variance. We assume that $V_{t}$ is positive-definite for each $t\geq 0$, meaning that $\langle\theta,V_{t}\theta\rangle\geq 0$ for all $\theta\in\mathbb{R}^{d}$. Typically, we will also have $V_{t}\preceq V_{t+1}$, where $\preceq$ is the positive semidefinite partial (Loewner) order.
We are interested obtaining bounds on $\|S_{\tau}\|_{V_{\tau}^{-1}}$ for any $\mathcal{F}$-adapted stopping time $\tau$ under various assumptions on the relationship between $S_{t}$ and $V_{t}$. More specifically, we think of $(V_{t})$ as the accumulated variance process of $(S_{t})$ or, in the words of Blackwell and Freedman [8], a measure of the intrinsic time-scale of $(S_{t})$. This is formalized with the notion of a sub-$\psi$ process, which generalizes many distributions of interest. Let $\mathbb{S}^{d-1}:=\{x\in\mathbb{R}^{d}:\|x\|_{2}=1\}$ denote the unit sphere in $\mathbb{R}^{d}$. Recall that a process $(A_{t})$ is a supermartingale with respect to $\mathcal{F}$ if it is $\mathcal{F}$-adapted and $\mathbb{E}[A_{t}|\mathcal{F}_{t-1}]\leq A_{t-1}$ for all $t\geq 1$.
@@
Let $\psi:[0,\lambda_{\max})\to\mathbb{R}_{\geq 0}$. Let $(S_{t})$ be an $\mathbb{R}^{d}$-valued process and $(V_{t})$, $V_{t}\succ 0$ an $\mathbb{R}^{d\times d}$-valued process, both adapted to a filtration $\mathcal{F}$. We say that $(S_{t})$ is a sub-$\psi$ process with variance proxy $(V_{t})$ if, for all $\theta\in\mathbb{S}^{d-1}$, all $\lambda\in[0,\lambda_{\max})$, and all $t\geq 0$,
@@
where $(L^{\lambda}_{t}(\theta))_{t\geq 0}$ is a nonnegative supermartingale adapted to $(\mathcal{F}_{t})$. Equivalently, we say that $(S_{t},V_{t})$ is a sub-$\psi$ process.
For scalar $S_{t}$ and $V_{t}$, sub-$\psi$ processes were originally proposed and studied by Howard et al. [34] for scalar and matrix settings, generalizing and unifying a wide array of earlier work, such as de la Peña et al. [17]. The vector case has been recently studied by Whitehouse et al. [63]. Like both Howard et al. [34] and Whitehouse et al. [63], we assume that $\psi$ is CGF-like, meaning that it is strictly convex and twice continuously differentiable, with $\lim_{\lambda\to 0^{+}}\textnormal{d}\psi(\lambda)/\textnormal{d}\lambda=0=\psi(0)$.
While the definition of a sub-$\psi$ process may appear abstract at first glance, let us reassure the reader that the definition captures many familiar (and unfamiliar) examples, including certain heavy-tailed settings.
@@
    $\psi_{N}=\psi_{N}(\lambda):=\lambda^{2}/2$ for $\lambda\in[0,\lambda_{\max})$ results in a sub-exponential process (sub-Gaussian if $\lambda_{\max}=\infty)$. If $X_{t}$ is $\Sigma_{t}$-sub-Gaussian [^1] That is, $\mathbb{E}[\exp\{\langle\theta,X_{t}\rangle)-\psi_{N}(\lambda)\langle\theta,\Sigma_{t}\theta\rangle\}|\mathcal{F}_{t-1}]\leq 1$ for all $\theta\in\mathbb{S}^{d-1}$ with $\lambda_{\max}=\infty$. In the scalar case, $X_{t}$ is $\sigma_{t}$-sub-Gaussian (conditional on the past) if $\mathbb{E}[\exp(\lambda X_{t})|\mathcal{F}_{t-1}]\leq\exp(\psi_{N}(\lambda)\sigma_{t}^{2})$ for all real $\lambda$. then $S_{t}=\sum_{k\leq t}X_{k}$ is a sub-$\psi_{N}$ process with $\lambda_{\max}=\infty$ and $V_{t}=\sum_{k\leq t}\Sigma_{k}$. Here $(\Sigma_{t})$ can be $\mathcal{F}$-adapted (not just predictable). As a heavy-tailed example, if $\mathbb{E}[\langle\theta,X_{t}\rangle|\mathcal{F}_{t-1}]<\infty$ for all $\theta\in\mathbb{S}^{d-1}$, then $S_{t}=\sum_{k\leq t}X_{k}$ is sub-$\psi_{N}$ with $\lambda_{\max}=\infty$ and $V_{t}=\frac{1}{3}\sum_{k\leq t}X_{k}X_{k}^{\intercal}+\frac{2}{3}\sum_{k\leq t}\mathbb{E}[X_{k}X_{k}^{\intercal}|\mathcal{F}_{k-1}]$ (see Howard et al. [34, Lemma 3f]). More exotically, consider any sequence of random vectors $(X_{t})$ that are conditionally symmetric: $X_{t}\sim-X_{t}|\mathcal{F}_{t-1}$ for all $t$. Lemma 3 of de la Peña et al. [19] implies that $S_{t}=\sum_{k\leq t}X_{k}$ and $V_{t}=\sum_{k\leq t}X_{k}X_{k}^{\intercal}$ define a sub-$\psi_{N}$ process with $\lambda_{\max}=\infty$; note here that we did not require even a single moment to exist, let alone a moment generating function, in order to yield a sub-$\psi_{N}$ process. As in the scalar case, bounded random vectors can also be shown to be sub-$\psi_{N}$.
@@
    $\psi_{G,c}(\lambda)=\frac{\lambda^{2}}{2(1-c\lambda)}$ for $c\in\mathbb{R}$ and $\lambda_{\max}=1/\max\{c,0\}$ results in a sub-gamma process; the terminology stems from the fact that $\psi_{G,c}$ is an upper bound on the CGF of a gamma random variable [10, Section 2.4]. As in the scalar case, random vectors whose moments in each direction obey a Bernstein-type inequality can be shown to be sub-$\psi_{G,c}$; see Lemma 5.2. Whitehouse et al. [63] further show that if $\mathbb{E}|\theta,X_{t}|^{3}$ is finite for all $\theta\in\mathbb{S}^{d-1}$, then $S_{t}=\sum_{j\leq t}X_{j}$ is sub-$\psi_{G,c}$ for $c=1/6$ and $V_{t}=\sum_{j\leq t}(X_{j}X_{j}^{\intercal}+\mathbb{E}[\|X_{j}\|^{3}|\mathcal{F}_{j-1}]I_{d})$. Moreover, Howard et al. [34, Proposition 1] show that any twice differentiable, CGF-like $\psi$ can be bounded by $a\psi_{G,c}$ for some $a,c>0$. Thus, any sub-$\psi$ process is sub-$\psi_{G,c}$ after scaling (though transforming to sub-gamma may cause some looseness in the resulting bounds).
@@
    $\psi_{E,c}(\lambda)=(-\log(1-c\lambda)-c\lambda)/c^{2}$ for $c\in\mathbb{R}$ where $\lambda_{\max}=1/\max\{c,0\}$ results in a sub-exponential process, which is equivalent [^2] In particular, Appendix E in Howard et al. [34] shows that $\log\mathbb{E}e^{\lambda X}\leq\psi_{E,c}(\lambda)\sigma^{2}$ for all $0\leq\lambda<1/c$ and some $\sigma,c>0$ if and only if $\log\mathbb{E}e^{\lambda X}\leq\psi_{N}(\lambda)u$ for all $0\leq\lambda<1/b$ for some $u,b>0$. up to constants to the sub-exponential condition mentioned under $\psi_{N}$; the terminology stems from the fact that $\psi_{E,c}$ is the CGF of a centered unit-rate negative-exponential random variable. Importantly, sums of bounded random vectors can be shown to be sub-exponential with an adapted (i.e. not $\mathcal{F}$-predictable) variance $V_{t}$; see Section 6. Indeed, this observation undergirds our empirical Bernstein bound. As in the sub-gamma case, Howard et al. [34, Proposition 1] also shows that any CGF-like $\psi$ can be upper bounded by $a\psi_{E,c}$ for some $a,c\geq 0$ (but as mentioned above for sub-gamma, exploiting this transformation may yield looser bounds than directly dealing with $\psi$).
@@
    $\psi_{P,c}(\lambda)=(e^{c\lambda}-c\lambda-1)/c^{2}$ for $c\in\mathbb{R}$ and $\lambda_{\max}=\infty$ results in a sub-Poisson process; this terminology stems from the fact that $\psi_{P,c}$ is the CGF of a centered unit-rate Poisson random variable. Most proofs of Bennett’s inequality (e.g., Boucheron et al. [10, Theorem 2.9]) involve showing that a bounded random variable is sub-Poisson. In the multivariate setting, Lemma 5.1 shows that vectors $(X_{t})$ with $\|X_{t}\|\leq b$ are sub-$\psi_{P,b}$ with $V_{t}=\sum_{j\leq t}\mathbb{E}[X_{j}X_{j}^{\intercal}|\mathcal{F}_{j-1}]$.
@@
To prove our results, we instead turn to a related but distinct technique which we call the variational approach to concentration. Here we also begin with a mixture but after step (2) applies the Donsker-Varadhan change of measure formula in order to bound $Z_{t}$ in terms of the KL divergence between $\rho$ and some prior distribution. The benefit of this technique is that $\rho$ can be data-dependent. This approach lies at the heart of (and originated in) PAC-Bayesian learning theory [31, 5]. A generic template for such a bound, which applies to general stochastic processes and handles stopping times, is given by Chugg et al. [15, Theorem 4]. See also Flynn et al. [25] for a similar statement.